# Simulation Based Bayesian Optimization

### v1

@@ 4.2 Contamination Control Problem
The contamination control problem was proposed by Hu et al. (2010). Consider a food supply with $d=25$ stages that may be contaminated with pathogenic microorganisms. $Z_{i}$ denotes the fraction of food contaminated at the $i$-th stage of the supply, for $1\leq i\leq d$. $Z$ evolves according to the random process described as follows. At stage $i$, a prevention effort (with fixed cost $c_{i}$) can be made that decreases contamination a random rate $\Gamma_{i}\in(0,1)$. If no prevention is taken, contamination spreads at a random rate $\Lambda_{i}\in(0,1)$. This results in the following recursive equation
@@
where $x_{i}\in\{0,1\}$ is the decision variable associated with the prevention effort at stage $i$, and $x_{i}=1$ denotes the the prevention effort was taken. The goal is to decide, at each stage $i$, whether to implement a prevention effort. The problem is constrained in that the cost should be minimized while ensuring that at each stage, the fraction of contaminated food does not exceed an upper limit $U_{i}$ with probability at least $1-\epsilon$. The random variables $\Gamma_{i}$ and $\Lambda_{i}$ as well as the initial fraction of contaminated food $Z_{1}$ follow beta distributions. We set $U_{i}=0.1$ for all $i$ and $\epsilon=0.05$, thus adopting the same framework as in Baptista and Poloczek (2018).
We consider the Lagrangian relaxation of the problem that is given by
@@
where each violation is penalized by $\rho=1$. We set $T=100$. An $l_{1}$ regularization term with parameter $\lambda=0.0001$ was added to encourage the prevention efforts to occur at a small number of stages.
Figure 2: Contamination control problem: average objective function values by number of evaluations, colored by algorithm. Small objective values indicate better performance.
@@ 4.3 RNA design
An RNA sequence can be represented as a string $A=a_{1}\cdots a_{p}$ of length $p$, where each $a_{i}\in\{A,U,G,C\}$ corresponds to a nucleotide letter. The secondary structure of this RNA is defined as the arrangement of base pairs that minimizes its free energy. Several RNA folding algorithms, such as the one proposed by Zuker and Stiegler (1981), employ dynamic programming and a thermodynamic model to estimate the Minimum Free Energy (MFE) of a given sequence. However, the computational complexity of these algorithms is $\mathcal{O}(p^{3})$, which becomes a bottleneck when dealing with a large number of RNA sequences. To address this challenge, we frame the RNA sequence optimization problem as in Dadkhahi et al. (2022): given a sequence of length $p$, the objective is to identify the RNA sequence that forms a secondary structure with the lowest MFE. In our optimization process, we utilize the well-known RNAfold package developed by Lorenz et al. (2011) to assess the MFE for a given RNA sequence. Our ultimate goal is to discover the RNA sequence with the lowest MFE while minimizing the number of calls made to the MFE evaluator. The results that follow correspond to sequences of length $p=30$.
Figure 3: RNA problem: average objective function values by number of evaluations, colored by algorithm. Small objective values of minimum free energy (MFE) indicate better performance.
@@ 4.4 General Comments
While most Bayesian Optimization literature have utilized Gaussian Processes as surrogate probabilistic models, our empirical study suggests that, for combinatorial optimization problems, alternative Bayesian models may lead to superior performance. Notably, in two out of three experiments, Sparse Bayesian linear regression with pairwise interactions demonstrated clear superiority over GPs and maintained competitiveness in the remaining experiment. This underscores the significance of exploring a broader range of Bayesian models within the realm of Bayesian Optimization.
@@ 5 Conclusion and Future Work
We have proposed SBBO, a simulation-based framework for optimizing acquisition functions in Bayesian Optimization that just requires sampling access to posterior predictive distributions of surrogate models. This flexibility enables the utilization of any Bayesian model whose posterior inference relies on MCMC as surrogate probabilistic model in BO. This can have advantages in non-standard black-box function optimization settings, as in the case of combinatorial search spaces. We have demonstrated empirically the effectiveness of SBBO under several surrogate probabilistic models in three combinatorial optimization problems.
Several future lines of research are possible. From a methodological perspective, we have only considered the greedy one-step look-ahead BO. SBBO could be extended to multi-step look-ahead BO. However, this extension is not direct. From a computational perspective we would suffer from the standard inefficiencies of dynamic programming. Some ideas in Bayesian sequential design might be relevant Wathen and Christen (2006); Brockwell and Kadane (2003); Drovandi et al. (2014). In terms of sampling strategies, we have proposed Metropolis-Hastings and Gibbs samplers to sample from the augmented distribution. Other options could be adopted to better explore high-dimensional multimodal surfaces, such as particle methods Amzal et al. (2006). In addition, when dealing with continuous search spaces, MCMC samplers that leverage gradient information, such as Hamiltonian Monte Carlo or the Metropolis-adjusted Langevin algorithm could be used.
On the applications side, given that Bayesian linear regression with pairwise interactions seems to work well, a model that captures interactions between three or more covariates is of interest. However, explicitly accounting for beyond-pairwise interactions becomes unfeasible when the dimensionality of the search space increases. As an alternative, some non-parametric approaches such as Bayesian additive regression tress Chipman et al. (2010) that capture relevant interactions between more than two covariates could be utilized. Finally, SBBO could be used with likelihood-free surrogate models, from which posterior predictive samples could be obtained via Approximate Bayesian Computation.
Acknowledgements: The authors gratefully acknowledge Spain’s consul in Chicago for his valuable help facilitating this collaboration during the pandemic. This project is based upon work supported by the Spanish state investigation agency under the Proyectos de Generación de Conocimiento 2022 grant No. PID2022-137331OB-C33.
Disclosure statement: The authors report there are no competing interests to declare.
SUPPLEMENTARY MATERIAL
@@ Appendix A Empirical assessment of convergence
We include Figure 4 to provide empirical guarantees of SBBO convergence. To make this plot, we initialized the contamination problem described in Section 4.2 with a dataset $\mathcal{D}_{1:100}$ of 100 of pairs of covariates and outcomes. Metropolis Hastings SBBO (Algorithm 1) with Tanimoto Gaussian Process regression as surrogate probabilistic model was used to find the subsequent evaluation location $x_{101}$. This process was repeated 40 times to account for the stochasticity of the SBBO algorithm. In each repetition, the cooling schedule incrementally increases $H$ from 50 to 2500 in steps of 10. For each $H$, we approximate the posterior predictive utility of the solution found by SBBO for that value of $H$ via Monte Carlo. Expected improvement was used as utility function. $H$ is represented in the x axis and the corresponding posterior predictive utility in the y axis. Each gray line corresponds to a repetition and the red line indicates the average.
Figure 4: SBBO with Metropolis-Hastings convergence

### v2

@@ 4.2 Contamination Control Problem
The contamination control problem was proposed by [14]. Consider a food supply with $d=25$ stages that may be contaminated with pathogenic microorganisms. $Z_{i}$ denotes the fraction of food contaminated at the $i$-th stage of the supply, for $1\leq i\leq d$. $Z$ evolves according to the following random process. At stage $i$, a prevention effort (with fixed cost $c_{i}$) can be made that decreases contamination a random rate $\Gamma_{i}\in(0,1)$. If no prevention is taken, contamination spreads at a random rate $\Lambda_{i}\in(0,1)$. This results in the recursive equation
@@
where $x_{i}\in\{0,1\}$ is the decision variable associated with the prevention effort at stage $i$, and $x_{i}=1$ denotes the the prevention effort was taken. The goal is to decide, at each stage $i$, whether to implement a prevention effort. The problem is constrained in that the cost should be minimized while ensuring that at each stage, the fraction of contaminated food does not exceed an upper limit $U_{i}$ with probability at least $1-\epsilon$. The random variables $\Gamma_{i}$ and $\Lambda_{i}$ as well as the initial fraction of contaminated food $Z_{1}$ follow beta distributions. We set $U_{i}=0.1$ for all $i$ and $\epsilon=0.05$, thus adopting the same framework as in [2].
We consider the Lagrangian relaxation of the problem given by
@@
where each violation is penalized by $\rho=1$. We set $T=100$. An $l_{1}$ regularization term with parameter $\lambda=0.0001$ is added to encourage the prevention efforts to occur at a small number of stages.
Figure 2: Contamination control problem: average $\pm$ one standard error of objective function values by number of function evaluations. Small objective values indicate better performance.
@@ 4.3 RNA design
An RNA sequence is represented as a string $A=a_{1}\cdots a_{p}$ of length $p$, where each $a_{i}\in\{A,U,G,C\}$ represents a nucleotide. The secondary structure of the RNA is defined by the arrangement of base pairs that minimizes its free energy, referred to as the Minimum Free Energy (MFE). The objective of RNA sequence optimization is to find the sequence that folds into the structure with the lowest MFE.
Several RNA folding algorithms, such as the one introduced by [30], use dynamic programming and thermodynamic models to estimate the MFE of a sequence. However, these algorithms have a computational complexity of $\mathcal{O}(p^{3})$, which makes them inefficient for evaluating a large number of sequences. The RNA sequence optimization problem becomes even more challenging as the number of possible sequences grows exponentially with $4^{p}$, making exhaustive search for the global minimum MFE impractical.
To overcome these challenges, we use BO to efficiently search for the optimal RNA sequence, reducing the number of MFE evaluations required. We use the RNAfold package [19] to compute the MFE for a given sequence. The results presented focus on sequences of length $p=30$.
Figure 3: RNA problem: average $\pm$ one standard error of objective function values by number of function evaluations. Small objective values of MFE indicate better performance.
@@ 4.4 General Comments
While most Bayesian Optimization literature have utilized Gaussian Processes as surrogate probabilistic models, our empirical study suggests that, for combinatorial optimization problems, alternative models may lead to superior performance. Notably, in two out of three experiments, Sparse Bayesian linear regression with pairwise interactions demonstrated clear superiority over GPs and maintained competitiveness in the remaining experiment. This underscores the significance of exploring a broader range of Bayesian models within the realm of Bayesian Optimization.
@@ 5 Conclusion and Future Work
We have proposed SBBO, a simulation-based framework for optimizing acquisition functions in Bayesian Optimization that just requires sampling access to posterior predictive distributions of surrogate models. This flexibility enables the utilization of any Bayesian model whose posterior inference relies on MCMC as surrogate probabilistic model in BO. This has advantages in non-standard black-box function optimization settings, specially in combinatorial search spaces. We have demonstrated empirically the effectiveness of SBBO under several surrogate probabilistic models in three combinatorial optimization problems.
Several future lines of research are possible. From a methodological perspective, we have only considered the greedy one-step look-ahead BO. SBBO could be extended to multi-step look-ahead BO. However, this extension is not direct. From a computational perspective we would suffer from the standard inefficiencies of dynamic programming. Some ideas in Bayesian sequential design might be relevant [28, 5, 7]. In terms of sampling strategies, we have proposed Metropolis-Hastings and Gibbs samplers to sample from the augmented distribution. Other options could be adopted to better explore high-dimensional multimodal surfaces, such as particle methods [1]. In addition, when dealing with continuous search spaces, MCMC samplers that leverage gradient information, such as Hamiltonian Monte Carlo or the Metropolis-adjusted Langevin algorithm could be used.
On the applications side, given that Bayesian linear regression with pairwise interactions seems to work well, a model that captures interactions between three or more covariates is of interest. However, explicitly accounting for beyond-pairwise interactions becomes unfeasible when the dimensionality of the search space increases. As an alternative, some non-parametric approaches such as Bayesian additive regression tress [6] that capture relevant interactions between more than two covariates could be utilized. Finally, SBBO could be used with likelihood-free surrogate models, from which posterior predictive samples could be obtained via Approximate Bayesian Computation.
@@ Appendix A Empirical assessment of convergence
We include Figure 4 to suggest empirical guarantees of SBBO convergence. To make this plot, we initialized the contamination problem described in Section 4.2 with a dataset $\mathcal{D}_{1:100}$ of 100 of pairs of covariates and outcomes. Metropolis Hastings SBBO (Algorithm 1) with Tanimoto Gaussian Process regression as surrogate probabilistic model was used to find the subsequent evaluation location $x_{101}$. This process was repeated 40 times to account for the stochasticity of the SBBO algorithm. At each repetition, the cooling schedule incrementally increases $H$ from 50 to 2500 in steps of 10. For each $H$, we approximate the posterior predictive utility of the solution found by SBBO for that value of $H$ via Monte Carlo. Expected improvement was used as utility function. $H$ is represented in the $x$-axis and the corresponding posterior predictive utility in the $y$-axis. Each gray line corresponds to a repetition and the red line indicates the average.
Figure 4: Posterior predictive utility versus $H$ for SBBO with Metropolis-Hastings.